OpenStudy (anonymous):

Solve each equation. Be sure to show all work in a neat and organized manner. 5) t – 13 = -29 6) 17 + d = -7 7) d+ 112 = 159 8) m/5 = -15 9) 0.4x = 16 10) x/12- = -8

OpenStudy (anonymous):

show all work? im afraid that we need you to show your work and we will help you

OpenStudy (anonymous):

I can't give you direct answers. However I hope this helps. t = -29 + 13 d = -7 - 17 m = -15 * 5 x = 16 /0.4 x = -8 * 12

OpenStudy (anonymous):

Thanks jazy

OpenStudy (anonymous):

Welcome (: